OpenStudy (faiqraees):

If \(\large \rm I_n = \int_0^1 x(1-x^3)^n dx\) Prove that \[\large \rm (3n+2)I_n = 3nI_{n-1}\]

OpenStudy (faiqraees):

@ganeshie8

OpenStudy (faiqraees):

If someone can post the solution or the starting integral form, it will be fine. I will work the rest out myself.

OpenStudy (faiqraees):

Nope it didn't help. It is calculus. The reference you gave was for mechanics.

OpenStudy (danni_joanaveil2):

im sorry hun i dont know then :(

OpenStudy (faiqraees):

No problem

OpenStudy (irishboy123):

\(I_n = \int\limits_0^1 x (1-x^3)^n dx\) \(I_n = \int\limits_0^1 x (1-x^3) (1-x^3)^{n-1} dx\) \(I_n = \int\limits_0^1 x (1-x^3)^{n-1} - x^4 (1-x^3)^{n-1} dx\) \(I_n = I_{n-1} - \int\limits_0^1 x^4 (1-x^3)^{n-1} dx\) \(I_n = I_{n-1} - \int\limits_0^1 x^2 .( x^2 (1-x^3)^{n-1}) dx\) IBP it now as per the hint i haven't actually done it but i think that should work

OpenStudy (faiqraees):

@Irishboy123 thank you very much

OpenStudy (kainui):

Fun looking problem

OpenStudy (faiqraees):

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Kainui Fun looking problem \(\color{#0cbb34}{\text{End of Quote}}\) Only mathematicians can consider a problem fun.